# Aouad, Ji, and Shaposhnik

### v1

@@ 1 Introduction
To put this model in concrete terms, consider job recruitment, an application discussed in Weitzman, (1979). Recent technological advances transform job recruitment processes in various industries. For example, video conferencing is often used for remote (offsite) interviews. More recently, AI systems have been introduced to conduct such initial remote interviews without human intervention; the transcriptions are then reviewed by human evaluators. Remote interviews provide convenient ways for companies to learn more about candidates at reduced costs, compared with onsite interviews. Nevertheless, remote interviews cannot fully replace onsite visits, which are often still necessary to eventually make job offers. Moreover, while cheaper than onsite visits, substantive remote interviews – beyond the preliminary screenings based on CVs – still require time from human interviewers or evaluators to assess the candidates’ fit for the position. In selecting a candidate for a vacant position, companies must therefore decide how to synthesize the different mediums to identify and select qualified candidates while minimizing costs. Put into the framework of our model, Pandora represents the company, boxes correspond to candidates whose prior evaluations are based on screenings and automated assessments, and box opening actions correspond to conducting interviews – offsite or onsite. The goal is to effectively administer the hiring process by deciding which candidates should be interviewed and in which format: cheaper and less accurate offsite interviews, or more expensive and informative onsite visits. While job recruitment serves as our running example, later in the paper, we will also draw connections to financial options.
@@ 1.1 Results and contributions
Hardness. We prove that computing optimal policies for PSI is NP-hard, contrary to the standard Pandora’s box problem. The proof is based on a reduction defined using PSI instances that consiss of boxes with binary prizes and where partial opening indicates one of two types.
Characterization of an optimal policy. Despite the hardness result, we identify threshold-based rules that almost fully describe the optimal policies. Similarly to the Pandora’s box problem, one type of threshold is associated with decisions to fully open boxes. To capture the added value of partial inspections, we define two additional types of thresholds associated with decisions to partially open boxes, which interestingly enough have an economic interpretation of strike prices of compound options. We show that, in most cases, the optimal actions can be prescribed using the newly defined thresholds. In particular, we characterize the optimal stopping criterion (i.e., when to stop searching and collect the best prize presently available), and identify sufficient conditions under which it is optimal to fully or partially open a box. We identify special non-trivial cases where our characterization fully describes an optimal policy, and show that a simple policy, which is defined using these thresholds, is asymptotically optimal. Numerical experiments show that optimal actions are identified by our theoretical characterization in more than 92% of all systems states, on average, and this percentage increases in the number of boxes of the problem instance.
Simple robust policies. Next, we focus on the class of committing policies.22endnote: 2Our approach is inspired by the class of committing algorithms studied by Beyhaghi and Kleinberg, (2019), with the notable difference that our setting involves sequential inspections. These semi-adaptive policies specify a priori whether each box should be partially or fully opened the first time it is inspected. Interestingly, each committing policy induces a corresponding multi-armed bandit problem (MAB) in which the optimal indices coincide with our proposed thresholds. By leveraging fundamental results on the adaptivity gap of stochastic submodular optimization problems, we argue that the optimal a committing policy can be efficiently approximated within a ratio of $1-1/e$.
@@
  • We examine whether there exists an optimal policy that can be expressed using index rules. We find that while, in general, the performance of index policies improves with the problem size, their capacity is limited, and such policies cannot fully describe optimal decisions throughout the entire state space. We introduce a generalized policy that more accurately represents optimal policies.
@@ 1.2 Related Work
The Pandora’s box problem and its variants. The seminal work of Weitzman, (1979) led to substantial research on more general and realistic settings. For example, Vishwanath, (1992) consider a generalization with parallel searches, where Pandora can open several boxes simultaneously at a discounted cost. The authors identify distributional assumptions under which it is optimal to open boxes according to Weitzman’s thresholds. Olszewski and Weber, (2015) examine a generalization where the objective function depends on all discovered prizes.44endnote: 4In the standard Pandora’s box problem, the objective function is simply the maximum observed prize. They show that if a threshold-based rule is unconditionally optimal, then the problem is reducible to the standard Pandora’s box setting. Boodaghians et al., (2020) study a variant of the Pandora’s box problem in which precedence constraints are imposed on the inspections of boxes. They show that when the precedence constraints are tree-shaped, the optimal policy can still be described as a threshold-based rule. Gibbard, (2022) extends the Pandora’s box by assuming that the prize is additive and that a two-step sequential inspection reveals the exact value of each part. This sequential inspection problem does not capture the tradeoff between cheap and costly inspections, which is the main focus of our work.
@@ 2 Problem Formulation
Pandora (the decision-maker) is given $N\in\mathbb{N}^{*}\textsection$ boxes containing random prizes. Each box $i\in[N]$ is associated with a priori unknown prize $V_{i}$ and type $T_{i}$ drawn from a known joint probability distribution $\mathcal{D}_{i}$. We assume that $V_{i}$ is either a continuous or discrete scalar random variable and that $T_{i}$ is a discrete scalar random variable with support $\Gamma_{i}$.66endnote: 6Throughout, we use capital letters to denote random variables and lowercase letters to denote their realizations. We assume that the support of $\mathcal{D}_{i}$ is contained in +×$\mathbb{R}^{+}\times\Gamma_{i}$, and that the random variables $(V_{1},T_{1}),\ldots,(V_{N},T_{N})$ are mutually independent (for each box $i$, the random variables $V_{i}$ and $T_{i}$ are dependent; otherwise the type would not be informative about the prize). Pandora can open a box in one of two inspection modes: full opening (F-opening) and partial opening (P-opening). The realization of type $T_{i}$ is revealed immediately by P-opening box $i\in[N]$, while the corresponding F-opening action reveals the realization of $(V_{i},T_{i})$. These opening actions incur constant costs, respectively denoted by $c_{i}^{F}$ and $c_{i}^{P}$, which are also known to Pandora.
@@ 2.2 Illustrative examples
There is a large number of i.i.d. boxes; each contains a prize of 110 with probability $\frac{1}{2}$ and zero otherwise. The box-opening costs are $c_{i}^{F}=100$ and $c_{i}^{P}=\epsilon$, and P-opening acquires the same information as F-opening (nevertheless, F-opening is still required to collect the prize). It is easy to see that given only the F-opening inspection mode, it is not worthwhile to F-open any box and the expected profit of any policy that opens boxes is negative. However, given the option to P-open, one could identify a box that contains the prize prior to F-opening it. In that case, the expected profit is $110-100=10$ minus $\epsilon$ times the expected number of opened boxes, which is equal to two.
@@
Consider the state ($(\{1\},\emptyset,y)$ where there is one box left with two types: ‘exceptional’ and ‘average,’ which can be revealed by P-opening. Box 1 is ‘exceptional’ with probability $\delta=0.01$ and is ‘average’ with probability $1-\delta$. If the box is exceptional, the prize is known to be $2$ almost surely, but if it is normal, the prize is uniformly distributed in $[0,1]$. Assuming that the F-opening cost is $c_{1}^{F}=\frac{1}{3}$ and the P-opening cost is $c_{1}^{P}=\frac{1}{3}\delta$, it can be easily seen that when $y=0$, both inspection modes are preferred over stopping (which yields 00). Moreover, F-opening dominates P-opening because F-opening is inevitable in this case. However, when $y=1$, F-opening would be no longer worthwhile compared to stopping, because the cost $c_{1}^{F}=\frac{1}{3}$ exceeds the expected gain $\delta\cdot(2-y)=\delta$. By contrast, P-opening (followed by the F-opening) of an exceptional box is advantageous because the expected cost is $c_{1}^{P}+\delta c_{1}^{F}=\frac{1}{3}\delta+\frac{1}{3}\delta=\frac{2}{3}\delta$, whereas the expected gain is $\delta$.
@@ 3.1 Opening thresholds and optimal stopping rule
The set of F-thresholds $\{\sigma_{i}^{F}\}$ in Equation (6) describes the optimal index policy for the standard Pandora’s box problem (Weitzman, 1979). The conventional economic interpretation for this opening threshold is that of a fair strike price of a hypothetical option contract. That is, if $c_{i}^{F}$ is the cost of a call option to buy prize $V_{i}$ with strike price $\sigma_{i}^{F}$, then $\sigma_{i}^{F}$ is the “fair” strike price that equalizes the cost and the expected profit of buying this option.
Much of this intuition carries over to our problem where inspections are sequential. Specifically, using the notion of compound option or, informally, a “two-stage call option.” The latter refers a more complex financial contract that allows the buyer to purchase an option at some future date. That is, for the price $c_{i}^{P}$ the decision-maker can buy the initial (overlying) option, which allows her at a future date (when more information is available) to purchase the second (underlying) option for a price of $c_{i}^{F}$. The additional information that is available at the time of purchasing the second option is represented by the realization $t_{i}$. The conditional F-threshold represents the fair strike price for purchasing $V_{i}$, given the information that is available at the future date, while the P-threshold represents the strike price for buying the initial option. This type of option is referred to as Call on Call. On the other hand, the FP-threshold corresponds to a compound option which is referred to as Call on Put. It allows the buyer to short an option at a future time. That is, in the first stage, the buyer pays $c_{i}^{P}$ to observe the type $t_{i}$, and then decides whether or not to enter the second stage. In the second stage, she has the right to choose whether to short a call option with prize $V_{i}$ and a strike price $\sigma_{i}^{F/P}$. Clearly, if the strike price $\sigma_{i}^{F/P}$ is overpriced, i.e., greater than $\sigma_{i}^{F\mid t_{i}}$, a risk-neutral player will sell the second-stage option and otherwise not. Therefore, we can interpret $\sigma_{i}^{F/P}$ as the fair strike price to equalize the expected costs and expected profits of this two-stage option.
The above interpretation motivates a policy in which the decision maker chooses the highest option strike price—that is, according to the box and opening mode with highest threshold. We show next that this index policy is indeed optimal for the single-box case, and we further expand its sufficient optimality conditions in the next sections.
@@
$-c_{i}^{F}+\mathbb{E}\Big{[}\max\left\{V_{i},y\right\}\Big{]}-y=-c_{i}^{F}+\mathbb{E}\Big{[}\max\{V_{i}-y,0\}\Big{]}=-\mathbb{E}\Big{[}\max\{V_{i}-\sigma_{i}^{F},0\}\Big{]}+\mathbb{E}\Big{[}\max\{V_{i}-y,0\}\Big{]}\ .$ (10)
Given the fair strike price interpretation of the F-threshold, there is a gain from F-opening if and only if $y\leq\sigma_{i}^{F}$, meaning the best prize at hand is lower than the fair strike price. Otherwise, it is preferable to stop searching. Similarly, the difference between the expected profit from P-opening and from stopping is
@@
Given the fair strike price interpretation of the P-threshold, there is a gain from P-opening if and only if $y\leq\sigma_{i}^{P}$, meaning the best prize at hand $y$ is lower than the fair strike price $\sigma_{i}^{P}$. Otherwise, it is preferable to stop searching.
@@
$=-c_{i}^{P}+\mathbb{E}_{T_{i}}\Big{[}\max\{0,c_{i}^{F}-\mathbb{E}_{V_{i}\mid T_{i}}\big{[}(V_{i}-y)^{+}\big{]}\}\Big{]}\ .$
With the fair strike price interpretation of the FP-threshold in mind, P-opening is better if and only if $y>\sigma^{F/P}_{i}$, meaning that we already have secured a prize $y$ higher than the fair strike price $\sigma^{F/P}_{i}$. Otherwise, it is preferable to F-open box $i$.
@@
$\forall i\in\mathcal{C}:y\geq\max\left\{\sigma_{i}^{F},\sigma_{i}^{P}\right\},\,\,\text{and }\,\,\forall j\in\mathcal{P}:y\geq\sigma_{j}^{F\mid t_{j}}.$ (12)
@@ 3.2 Sufficient conditions for optimal F-opening
Having characterized the optimal stopping criterion, we now focus on states $(\mathcal{C},\mathcal{P},y)$ in which stopping is not optimal. We define the largest opening threshold $\sigma_{M}(\mathcal{C},\mathcal{P})$:
Note that the threshold $y$ associated with stopping is not part of this expression as we assumed that condition (12) does not hold. We refer to the box attaining the largest threshold as the leading box. Here, there are two possibilities: $\sigma_{M}(\mathcal{C},\mathcal{P})$ corresponds either to an F-threshold or a P-threshold. The next theorem establishes that in the former case, it is optimal to F-open the leading box. This condition holds regardless of whether the leading box is closed or partially-opened.
@@ 3.4 An optimal policy for a problem with binary prizes
Next, we study a special case of PSIs whereby boxes are statistically identical and independent, contain binary Bernoulli prizes, and belong to one of two types—“good” or “bad.” We refer to this problem as PSI-B2I (Bernoulli, 2-type, identical and independent). This setting applies to situations where a decision maker is searching for a satisfactory alternative among options that are initially perceived as equivalent. Partial inspection reveals whether an option is promising (“good”) or has low potential (“bad”), while full inspection discloses the exact outcome. Several variants of this problem in which the decision maker cannot inspect boxes sequentially have been studied in the literature (see, e.g. Doval, 2018, and the references therein). Our setting is also different from Clarkson et al., (2020) who assume that there is exactly one prize hidden in the boxes and that opening can be done repeatedly (which does not fit our motivating applications).
@@
This above policy structure captures the case $\sigma^{P}\leq\sigma^{F}$, using the convention $N_{c}=\infty$. We also observe that the special case where $\sigma^{F}<\sigma^{P}$ and $\sigma^{F/P}\leq 0$ is subsumed by Theorem 3.5 (see Observation 1), in which case $N_{c}=0$. Therefore, to ease our analysis, we make without loss of generality the following simplifying assumption: {assumption} The boxes of PSI-B2I satisfy: $0<\sigma^{F/P}<\sigma^{F}<\sigma^{P}.$
@@
The proof can be found in Appendix 8.6.

### v2

# MOOR-0001-2024.00

@@ 1 Introduction
To put this model in concrete terms, consider job recruitment, an application discussed in Weitzman, (1979). Recent technological advances transform job recruitment processes in various industries. For example, video conferencing is often used for remote (offsite) interviews. More recently, AI systems have been introduced to conduct such initial remote interviews without human intervention; the transcriptions are then reviewed by human evaluators. Remote interviews provide convenient ways for companies to learn more about candidates at reduced costs, compared with onsite interviews. Nevertheless, remote interviews cannot fully replace onsite visits, which are often still necessary to eventually make job offers. Moreover, while cheaper than onsite visits, substantive remote interviews – beyond the preliminary screenings based on CVs – still require time from human interviewers or evaluators to assess the candidates’ fit for the position. In selecting a candidate for a vacant position, companies must therefore decide how to synthesize the different mediums to identify and select qualified candidates while minimizing costs. Put into the framework of our model, Pandora represents the company, boxes correspond to candidates whose prior evaluations are based on screenings and automated assessments, and box opening actions correspond to conducting interviews – offsite or onsite. The goal is to effectively administer the hiring process by deciding which candidates should be interviewed and in which format: cheaper and less accurate offsite interviews, or more expensive and informative onsite visits.
@@ 1.1 Results and contributions
Hardness. We prove that computing optimal policies for PSI is NP-hard, contrary to the standard Pandora’s box problem. The proof is based on a reduction defined using PSI instances that consist of boxes with binary prizes and where partial opening indicates one of two types.
Characterization of an optimal policy. Despite the hardness result, we identify threshold-based rules that almost fully describe the optimal policies. Similarly to the Pandora’s box problem, one type of threshold is associated with decisions to fully open boxes. To capture the added value of partial inspections, we define two additional types of thresholds associated with decisions to partially open boxes. We show that, in most cases, the optimal actions can be prescribed using the newly defined thresholds. In particular, we characterize the optimal stopping criterion (i.e., when to stop searching and collect the best prize presently available), and identify sufficient conditions under which it is optimal to fully or partially open a box. We identify special non-trivial cases where our characterization fully describes an optimal policy, and show that a simple policy, which is defined using these thresholds, is asymptotically optimal. Numerical experiments show that optimal actions are identified by our theoretical characterization in more than 92% of all systems states, on average, and this percentage increases in the number of boxes of the problem instance.
Simple robust policies. Next, we focus on the class of committing policies.22endnote: 2Our approach is inspired by the class of committing algorithms studied by Beyhaghi and Kleinberg, (2019), with the notable difference that our setting involves sequential inspections. These semi-adaptive policies specify a priori whether each box should be partially or fully opened the first time it is inspected. Interestingly, each committing policy induces a corresponding multi-armed bandit problem (MAB) in which the optimal indices coincide with our proposed thresholds. By leveraging fundamental results on the adaptivity gap of stochastic submodular optimization problems, we argue that the optimal a committing policy can be efficiently approximated within a ratio of 11/e1-1/e.
@@ 1.2 Related Work
The Pandora’s box problem and its variants. The seminal work of Weitzman, (1979) led to substantial research on more general and realistic settings. For example, Vishwanath, (1992) consider a generalization with parallel searches, where Pandora can open several boxes simultaneously at a discounted cost. The authors identify distributional assumptions under which it is optimal to open boxes according to Weitzman’s thresholds. Olszewski and Weber, (2015) examine a generalization where the objective function depends on all discovered prizes.44endnote: 4In the standard Pandora’s box problem, the objective function is simply the maximum observed prize. They show that if a threshold-based rule is unconditionally optimal, then the problem is reducible to the standard Pandora’s box setting. Boodaghians et al., (2020) study a variant of the Pandora’s box problem in which precedence constraints are imposed on the inspections of boxes. They show that when the precedence constraints are tree-shaped, the optimal policy can still be described as a threshold-based rule. Gibbard, (2022) extend the Pandora’s box problem by assuming that the prize is additive and that a two-step sequential inspection reveals the exact value of each part. Conceptually, such a model studies how information acquisition should be prioritized when the acquisition process can be broken into multiple steps. The model is a special case of the MAB problem (Bertsimas and Nino-Mora, 1996) that inherits the indexability structure of the optimal policy. In contrast, our model attempts to capture a different tradeoff between two information acquisition methods (a cheaper and less precise versus a more expensive and accurate), a characteristic which we show to be challenging analytically and computationally.
@@ 2 Problem Formulation
Pandora (the decision-maker) is given $N\in\mathbb{N}^{*}\textsection$ boxes containing random prizes. Each box $i\in[N]$ is associated with a priori unknown prize $V_{i}$ and type $T_{i}$ drawn from a known joint probability distribution $\mathcal{D}_{i}$. We assume that $V_{i}$ is either a continuous or discrete scalar random variable and that $T_{i}$ is a discrete scalar random variable with support $\Gamma_{i}$.66endnote: 6Throughout, we use capital letters to denote random variables and lowercase letters to denote their realizations. We assume that the support of $\mathcal{D}_{i}$ is contained in $\mathbb{R}^{+}\times\Gamma_{i}$, and that the random variables $(V_{1},T_{1}),\ldots,(V_{N},T_{N})$ are mutually independent (for each box $i$, the random variables $V_{i}$ and $T_{i}$ are dependent; otherwise the type would not be informative about the prize). Pandora can open a box in one of two inspection modes: full opening (F-opening) and partial opening (P-opening). The realization of type $T_{i}$ is revealed immediately by P-opening box $i\in[N]$, while the corresponding F-opening action reveals the realization of $(V_{i},T_{i})$. These opening actions incur constant costs, respectively denoted by $c_{i}^{F}$ and $c_{i}^{P}$, which are also known to Pandora.
@@ 2.2 Illustrative examples
There is a large number of i.i.d. boxes; each contains a prize of 110 with probability $\frac{1}{2}$ and zero otherwise. The box-opening costs are $c_{i}^{F}=100$ and $c_{i}^{P}=\epsilon$, and P-opening acquires the same information as F-opening (nevertheless, F-opening is still required to collect the prize). It is easy to see that given only the F-opening inspection mode, it is not worthwhile to F-open any box and the expected profit of any policy that opens boxes is negative. However, given the option to P-open, one could identify a box that contains the prize prior to F-opening it. In that case, the expected profit is 110100=10110-100=10 minus $\epsilon$ times the expected number of opened boxes, which is equal to two.
@@
Consider the state $(\{1\},\emptyset,y)$ where there is one box left with two types: ‘exceptional’ and ‘average,’ which can be revealed by P-opening. Box 1 is ‘exceptional’ with probability $\delta=0.01$ and is ‘average’ with probability $1-\delta$. If the box is exceptional, the prize is known to be $2$ almost surely, but if it is average, the prize is uniformly distributed in $[0,1]$. Assuming that the F-opening cost is $c_{1}^{F}=\frac{1}{3}$ and the P-opening cost is $c_{1}^{P}=\frac{1}{3}\delta$, it can be easily seen that when $y=0$, both inspection modes are preferred over stopping (which yields 0). Moreover, F-opening dominates P-opening because F-opening is inevitable in this case. However, when $y=1$, F-opening would be no longer worthwhile compared to stopping, because the cost $c_{1}^{F}=\frac{1}{3}$ exceeds the expected gain $\delta\cdot(2-y)=\delta$. By contrast, P-opening (followed by the F-opening) of an exceptional box is advantageous because the expected cost is $c_{1}^{P}+\delta c_{1}^{F}=\frac{1}{3}\delta+\frac{1}{3}\delta=\frac{2}{3}\delta$, whereas the expected gain is $\delta$.
@@ 3.1 Opening thresholds and optimal stopping rule
The set of F-thresholds $\{\sigma_{i}^{F}\}$ in Equation (6) describes the optimal index policy for the standard Pandora’s box problem (Weitzman, 1979). The conventional economic interpretation for this opening threshold is that of a fair strike price of a hypothetical option contract. That is, if $c_{i}^{F}$ is the cost of a call option to buy prize $V_{i}$ with strike price $\sigma_{i}^{F}$, then $\sigma_{i}^{F}$ is the “fair” strike price that equalizes the cost and the expected profit of buying this option. This intuition carries over to our problem where inspections are sequential using the notion of a compound option or, informally, a “two-stage call option.” This analogy is developed in Appendix 8.
@@
$c_{i}^{F}+\mathbb{E}\Big[\max\left\{V_{i},y\right\}\Big]-y=-c_{i}^{F}+\mathbb{E}\Big[\max\{V_{i}-y,0\}\Big]=-\mathbb{E}\Big[\max\{V_{i}-\sigma_{i}^{F},0\}\Big]+\mathbb{E}\Big[\max\{V_{i}-y,0\}\Big]\ .$ (10)
The F-threshold can be therefore interpreted as a break-even point for the value of $y$ under which F-opening is preferred to stopping, and over which stopping is preferred. Similarly, in this single-box case, the difference between the expected profit from P-opening (and following the optimal action thereafter) and from stopping is
@@
Similarly, the P-threshold can be interpreted as a break-even point for the value of $y$ under which P-opening is preferred to stopping, and over which stopping is preferred.
@@
$=-c_{i}^{P}+\mathbb{E}_{T_{i}}\Big[\max\{0,c_{i}^{F}-\mathbb{E}_{V_{i}\mid T_{i}}\big[(V_{i}-y)^{+}\big]\}\Big]\ .$ (12)
One can verify that the threshold $\sigma_{i}^{F/P}$ is a break-even point on the value of $y$ that determines whether P-opening or F-opening achieves a higher expected profit. Intuitively, in the single-box case, the definition for the set of thresholds $\sigma_{i}^{F}$, $\sigma_{i}^{F\mid t_{i}}$, $\sigma_{i}^{P}$, and $\sigma_{i}^{F/P}$ exploits the single-crossing property between pairs of action-value functions. In the more general case, the definition can be viewed as using a myopic approximation for the action-value functions that assumes that only one box may be opened.
@@
We define the largest opening threshold $\sigma_{M}(\mathcal{C},\mathcal{P})$:
We refer to the box attaining the largest threshold $\sigma_{M}(\mathcal{C},\mathcal{P})$ as the leading box.
@@
$y\geq\sigma_{M}(\mathcal{C},\mathcal{P}).$ (14)
@@ 3.2 Sufficient conditions for optimal F-opening
Having characterized the optimal stopping criterion, we now focus on states $(\mathcal{C},\mathcal{P},y)$ in which stopping is not optimal. Here, there are two possibilities: $\sigma_{M}(\mathcal{C},\mathcal{P})$ corresponds either to an F-threshold or a P-threshold. The next theorem establishes that in the former case, it is optimal to F-open the leading box. This condition holds regardless of whether the leading box is closed or partially-opened.
@@ 3.4 An optimal policy for a problem with binary prizes
Next, we study a special case of PSIs whereby boxes are statistically identical and independent, contain binary Bernoulli prizes, and belong to one of two types—“good” or “bad.” We refer to this problem as PSI-B2I (Bernoulli, 2-type, identical and independent). This setting applies to situations where a decision maker is searching for a satisfactory alternative among options that are initially perceived as equivalent. Partial inspection reveals whether an alternative is promising (“good”) or has low potential (“bad”), while full inspection discloses the exact outcome. Several variants of this problem in which the decision maker cannot inspect boxes sequentially have been studied in the literature (see, e.g. Doval, 2018, and the references therein). Our setting is also different from Clarkson et al., (2020) who assume that there is exactly one prize hidden in the boxes and that opening can be done repeatedly (which does not fit our motivating applications).
@@
The structure of the policy above captures the case $\sigma^{P}\leq\sigma^{F}$, using the convention $N_{c}=\infty$. We also observe that the special case where $\sigma^{F}<\sigma^{P}$ and $\sigma^{F/P}\leq 0$ is subsumed by Theorem 3.5 (see Observation 1), in which case $N_{c}=0$. Therefore, to ease our analysis, we make without loss of generality the following simplifying assumption: {assumption} The boxes of PSI-B2I satisfy: $0<\sigma^{F/P}<\sigma^{F}<\sigma^{P}.$
@@
The proof can be found in Appendix 9.6. The theorem is presented as an algorithm in Appendix 9.7.
@@
Observe that this policy satisfies the optimal characterization given in Theorems 3.2-3.5; indeed, the proof in Appendix 9.6 directly builds on these properties. Interestingly, although PSI-B2I is a special case, much of the structure exhibited by the optimal policy for this instance extends naturally to the broader class of PSI problems. For example, it is useful to consider a class of policies that commit ex ante to fully or partially opening each closed box before execution; this notion of a committing policy proves useful in the general problem and is discussed in detail in Section 4.3. As another illustration, the instance highlights how an increase in the number of boxes makes partial opening more advantageous, a phenomenon that we further validate numerically in Section 5.4.